Required information

The Foundational 15 [LO6-1, LO6-2, LO6-3, LO6-4, LO6-5]

[The following information applies to the questions displayed below.]

Diego Company manufactures one product that is sold for $70 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 53,000 units and sold 48,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 21
Direct labor $ 10
Variable manufacturing overhead $ 2
Variable selling and administrative $ 4
Fixed costs per year:
Fixed manufacturing overhead $ 1,060,000
Fixed selling and administrative expense $ 557,000

The company sold 36,000 units in the East region and 12,000 units in the West region. It determined that $270,000 of its fixed selling and administrative expense is traceable to the West region, $220,000 is traceable to the East region, and the remaining $67,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

9. If the sales volumes in the East and West regions had been reversed, what would be the company’s overall break-even point in unit sales?
Answer #1

Total fixed cost = 1060000+557000 = 1617000

variable cost per unit = 21+10+2+4 = 37 per unit

Contribution margin per unit = 70-37 = 33 per unit

Overall break even = Fixed cost/Contribution margin per unit = 1617000/33 = 49000 Units